# Infinite graph product of groups I: Geometry of the extension graph

### v2

@@ 4 Asymptotic dimension of the extension graph
We now prove Theorem 4.4 using Corollary 3.20 and the natural contraction $\Gamma^{e}\to\widehat{\Gamma}^{e}$. See Section 2.3 for relevant notions.
@@
When $R=0$, we have $f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,0))=\{o\}$ for any $o\in V(\Gamma^{e})$ and the family ($(\{o\})_{o\in V(\Gamma^{e})}$ satisfies $\mathrm{asdim}\leq 0\,(\leq n)$ uniformly. Hence, the claim holds for $R=0$.
@@
Let $r\in\mathbb{N}$. For each $o\in V(\Gamma^{e})$, define $Y_{r,o}$ by $Y_{r,o}=\mathcal{N}_{\Gamma^{e}}(f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R)),3r+12)\cap X_{o}$. Since the family $\big{(}\,f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R))\,\big{)}_{o\in V(\Gamma^{e})}$ satisfies $\mathrm{asdim}\leq n$ uniformly by our assumption of induction, it’s not difficult to see that the family $(Y_{r,o})_{o\in V(\Gamma^{e})}$ also satisfies $\mathrm{asdim}\leq n$ uniformly. Let $o\in V(\Gamma^{e})$. Suppose that $(k,g),(\ell,h)\in\{0,\cdots,R+1\}\times\Gamma\mathcal{G}$ satisfy $Q_{o}(k,g)\neq Q_{o}(\ell,h)$ and that we have $a\in Q_{o}(k,g)\setminus Y_{r,o}$ and $b\in Q_{o}(\ell,h)\setminus Y_{r,o}$. By $\{a,b\}\subset X_{o}\setminus Y_{r,o}$, we have $d_{\widehat{\Gamma}^{e}}(o,a)=d_{\widehat{\Gamma}^{e}}(o,b)=R+1$, hence $k=\ell=R+1$. By this and $Q_{o}(k,g)\neq Q_{o}(\ell,h)$, we have $g\neq h$. By $a,b\notin\mathcal{N}_{\Gamma^{e}}(f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R)),3r+12)$, we also have $\min\{d_{\Gamma^{e}}(a,p_{R}(o,a)),d_{\Gamma^{e}}(b,p_{R}(o,b))\}>3r+12$.
@@ 5.2 Tightness in the sense of Bowditch
$A_{1}$ =$=\{g\in A\mid d_{\Gamma^{e}}(p_{-},p_{out}(g.\Gamma))\leq d_{\Gamma^{e}}(p_{-},c)\},$
$A_{2}$ =$=\{g\in A\mid d_{\Gamma^{e}}(p_{-},c)\leq d_{\Gamma^{e}}(p_{-},p_{in}(g.\Gamma))\},$
$A_{3}$ =$=\{g\in A\mid c\in V(g.\Gamma\cap p)\}.$
@@
$V(a,b)$ =$=\bigcup\{V(p)\mid p\in\mathrm{Geo}_{\Gamma^{e}}(a,b)\},$
$V(a,b\,;r)$ =$=\bigcup\{V(a^{\prime},b^{\prime})\mid a^{\prime}\in\mathcal{N}_{\Gamma^{e}}(a,r),\,b^{\prime}\in\mathcal{N}_{\Gamma^{e}}(b,r)\}.$
@@
$A$ =$=\{e\in E(p)\mid d_{\Gamma^{e}}(c,e)\leq k+2\delta+2\},$
$B$ =$=\{g\in S(p\,;5)\mid d_{\Gamma^{e}}(c,g.\Gamma\cap p)\leq k+2\delta+2\},$
$C$ =$=\{v\in V(\Gamma^{e})\mid\exists\,e\in A,\exists\,g\in B{\rm~{}s.t.~{}}v\in C(e,24\delta+32,g)\}.$
@@
$A$ =$=\{e\in E(p)\mid d_{\Gamma^{e}}(c,e)\leq k+4\delta+2\},$
$B$ =$=\{g\in S(p\,;3)\mid d_{\Gamma^{e}}(c,g.\Gamma\cap p)\leq k+4\delta+2\},$
$C$ =$=\{v\in V(\Gamma^{e})\mid\exists\,e\in A,\exists\,g\in B{\rm~{}s.t.~{}}v\in C(e,32\delta+32,g)\}.$
@@ 5.3 Fineness
$L(e,n)$ =$=\bigcup\{E(\gamma)\mid\gamma\in\mathscr{C}_{\Gamma^{e}}(e,n)\},$
$P(e)$ =$=\{g\in\Gamma\mathcal{G}\mid e\subset g.\Gamma\},$
$M(e,n)$ =$=\bigcup\{E(\gamma)\mid\text{$\gamma\in\mathscr{C}_{\Gamma^{e}}(e,n)$ s.t. $\exists\,g\in P(e)$ with $\gamma\subset g.\Gamma$}\}.$
@@
6 Relative hyperbolicity of graph-wreath product
@@
is a group homomorphism. Hence, $\alpha$ defines the semi-direct product $\Gamma\mathcal{G}\rtimes G$, which is called graph-wreath product.
@@
    For any $v\in V(G)$, $\mathrm{Stab}_{G}(v)$ is finitely generated.

### v3

@@ 4 Asymptotic dimension of the extension graph
We now prove Theorem 4.5 using Corollary 3.20 and the natural contraction $\Gamma^{e}\to\widehat{\Gamma}^{e}$. See Section 2.3 for relevant notions. Lemma 4.4 is an auxiliary lemma for Theorem 4.5.
Lemma 4.4.
Let $(S_{\alpha},d_{\alpha})_{\alpha\in\mathscr{A}}$ be a family of metric spaces and $(A_{\alpha})_{\alpha\in\mathscr{A}}$ be a family of subsets of $(S_{\alpha})_{\alpha\in\mathscr{A}}$ (i.e. $\forall\,\alpha\in\mathscr{A},\,A_{\alpha}\subset S_{\alpha}$) satisfying $\mathrm{asdim}\leq n$ uniformly with $n\in\mathbb{N}\cup\{0\}$. Then, the following hold.
  • (1)
    For any $t\geq 0$, the family $\big(\mathcal{N}_{S_{\alpha}}(A_{\alpha},t)\big)_{\alpha\in\mathscr{A}}$ satisfies $\mathrm{asdim}\leq n$ uniformly.
  • (2)
    Any family $(B_{\alpha})_{\alpha\in\mathscr{A}}$ of subsets of $(A_{\alpha})_{\alpha\in\mathscr{A}}$ satisfies $\mathrm{asdim}\leq n$ uniformly.
Proof.
[…]
(2) Let $r>0$. Since $(A_{\alpha})_{\alpha\in\mathscr{A}}$ satisfies $\mathrm{asdim}\leq n$ uniformly, there exist $\mathcal{U}^{0}_{\alpha},\cdots,\mathcal{U}^{n}_{\alpha}\subset 2^{A_{\alpha}}$ for each $\alpha\in\mathscr{A}$ satisfying the three conditions of Definition 2.23 for $r$. The family $\mathcal{V}^{0}_{\alpha},\cdots,\mathcal{V}^{n}_{\alpha}\subset 2^{B_{\alpha}}$ defined by $\mathcal{V}^{i}_{\alpha}=\big\{U\cap B_{\alpha}\in 2^{B_{\alpha}}\mid U\in\mathcal{U}^{i}_{\alpha}\big\}$ for each $i$ and $\alpha$ satisfies the three conditions of Definition 2.23 for $r$. Hence, $(B_{\alpha})_{\alpha\in\mathscr{A}}$ satisfies $\mathrm{asdim}\leq n$ uniformly. ∎
@@
When $R=0$, we have $f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,0))=\{o\}$ for any $o\in V(\Gamma^{e})$ and the family $(\{o\})_{o\in V(\Gamma^{e})}$ satisfies $\mathrm{asdim}\leq 0\,(\leq n)$ uniformly. Hence, the claim holds for $R=0$.
@@
Let $r\in\mathbb{N}$. For each $o\in V(\Gamma^{e})$, define $Y_{r,o}$ by $Y_{r,o}=\mathcal{N}_{\Gamma^{e}}(f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R)),3r+12)\cap X_{o}$. Since the family $\big(\,f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R))\,\big)_{o\in V(\Gamma^{e})}$ satisfies $\mathrm{asdim}\leq n$ uniformly by our assumption of induction, the family $(Y_{r,o})_{o\in V(\Gamma^{e})}$ also satisfies $\mathrm{asdim}\leq n$ uniformly by Lemma 4.4. (Here, we first applied Lemma 4.4 (1) with $\mathscr{A}=V(\Gamma^{e})$, $S_{\alpha}\equiv\Gamma^{e}$, and $t=3r+12$, and then applied Lemma 4.4 (2) by using $Y_{r,o}\subset\mathcal{N}_{\Gamma^{e}}(f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R)),3r+12)$.)
Let $o\in V(\Gamma^{e})$. Suppose that $(k,g),(\ell,h)\in\{0,\cdots,R+1\}\times\Gamma\mathcal{G}$ satisfy $Q_{o}(k,g)\neq Q_{o}(\ell,h)$ and that we have $a\in Q_{o}(k,g)\setminus Y_{r,o}$ and $b\in Q_{o}(\ell,h)\setminus Y_{r,o}$. By $\{a,b\}\subset X_{o}\setminus Y_{r,o}$, we have $d_{\widehat{\Gamma}^{e}}(o,a)=d_{\widehat{\Gamma}^{e}}(o,b)=R+1$, hence $k=\ell=R+1$. By this and $Q_{o}(k,g)\neq Q_{o}(\ell,h)$, we have $g\neq h$. By $a,b\notin\mathcal{N}_{\Gamma^{e}}(f^{-1}(\mathcal{N}_{\widehat{\Gamma}^{e}}(o,R)),3r+12)$, we also have $\min\{d_{\Gamma^{e}}(a,p_{R}(o,a)),d_{\Gamma^{e}}(b,p_{R}(o,b))\}>3r+12$.
@@ 5.2 Tightness in the sense of Bowditch
$A_{1}$ $=\{g\in A\mid d_{\Gamma^{e}}(p_{-},p_{out}(g.\Gamma))\leq d_{\Gamma^{e}}(p_{-},c)\},$
$A_{2}$ $=\{g\in A\mid d_{\Gamma^{e}}(p_{-},c)\leq d_{\Gamma^{e}}(p_{-},p_{in}(g.\Gamma))\},$
$A_{3}$ $=\{g\in A\mid c\in V(g.\Gamma\cap p)\}.$
@@
$V(a,b)$ $=\bigcup\{V(p)\mid p\in\mathrm{Geo}_{\Gamma^{e}}(a,b)\},$
$V(a,b\,;r)$ $=\bigcup\{V(a^{\prime},b^{\prime})\mid a^{\prime}\in\mathcal{N}_{\Gamma^{e}}(a,r),\,b^{\prime}\in\mathcal{N}_{\Gamma^{e}}(b,r)\}.$
@@
$A$ $=\{e\in E(p)\mid d_{\Gamma^{e}}(c,e)\leq k+2\delta+2\},$
$B$ $=\{g\in S(p\,;5)\mid d_{\Gamma^{e}}(c,g.\Gamma\cap p)\leq k+2\delta+2\},$
$C$ $=\{v\in V(\Gamma^{e})\mid\exists\,e\in A,\exists\,g\in B{\rm~s.t.~}v\in C(e,24\delta+32,g)\}.$
@@
$A$ $=\{e\in E(p)\mid d_{\Gamma^{e}}(c,e)\leq k+4\delta+2\},$
$B$ $=\{g\in S(p\,;3)\mid d_{\Gamma^{e}}(c,g.\Gamma\cap p)\leq k+4\delta+2\},$
$C$ $=\{v\in V(\Gamma^{e})\mid\exists\,e\in A,\exists\,g\in B{\rm~s.t.~}v\in C(e,32\delta+32,g)\}.$
@@
When $\Gamma$ is finite, the statement of Corollary 5.6 essentially follows from [Val21, Corollary C] since the contact graph and crossing graph are quasi-isometric (see [Gen25, Proposition 3.3] and Proposition 3.19).
@@ 5.3 Fineness
$L(e,n)$ $=\bigcup\{E(\gamma)\mid\gamma\in\mathscr{C}_{\Gamma^{e}}(e,n)\},$
$P(e)$ $=\{g\in\Gamma\mathcal{G}\mid e\subset g.\Gamma\},$
$M(e,n)$ $=\bigcup\{E(\gamma)\mid\text{$\gamma\in\mathscr{C}_{\Gamma^{e}}(e,n)$ s.t. $\exists\,g\in P(e)$ with $\gamma\subset g.\Gamma$}\}.$
@@
Lemma 5.11.
Proof.
[…]
6 Relative hyperbolicity of graph wreath product
@@
is a group homomorphism. Hence, $\alpha$ defines the semi-direct product $\Gamma\mathcal{G}\rtimes G$, which is called graph wreath product.
@@
    For any $v\in V(\Gamma)$, $\mathrm{Stab}_{G}(v)$ is finitely generated.